OpenStudy (anonymous):

fan and medal for best answer :) an expression is shown below: f(x)=-16x^2+8x+3 part a.) what are the x intercepts of the graph of f(x)? show work part b.) is the vertex of the graph of f(x) going to be maximum or minimum? what are the coordinates of the vertex? justify answer and show work part c.) what are the steps you would use to graph f(x)? make sure you can use answer obtained in part a and part b to draw graph.

OpenStudy (precal):

just sub 0 for f(x) and solve for x. I would use the quadratic formula for part a

OpenStudy (precal):

part b is a maximum and we know that because of the -16x^2 that means it is concave down

OpenStudy (precal):

use x=-b/2a to find the vertex location

OpenStudy (precal):

then take that value and sub it into the original to find the y value of the vertex

OpenStudy (anonymous):

i don't know how to use the quadratic formula

OpenStudy (precal):

I would graph using the vertex, y intercept and x intercepts

OpenStudy (precal):

http://www.purplemath.com/modules/quadform.htm

OpenStudy (precal):

if you are in Algebra I, do not bypass the quadratic formula. You will need it throughout your entire math studies.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how do you get the vertex coordinates @precal i don't get it

OpenStudy (precal):

do you have a graphing calculator so that you can verify the graph and the solutions

OpenStudy (anonymous):

coordinate of the vertex is ( 1/4, 4)

OpenStudy (anonymous):

how would you answer part c?

OpenStudy (precal):

let me check it with my graphing calculator

OpenStudy (precal):

not exactly

OpenStudy (precal):

a=-16 b=8 and c = 3

OpenStudy (anonymous):

what?????

OpenStudy (precal):

x=-b/2a x=-(8)/(2*-16) x=-8/-32 x=8/32 x=1/4 sorry you were correct, my calculator gave me x=.2500078

OpenStudy (precal):

I just showed you how to find the x coordinate of your vertex, now sub x=1/4 into the function and solve for your y value

OpenStudy (anonymous):

confused :/

OpenStudy (precal):

Are you in Algebra 1?

OpenStudy (anonymous):

yes

OpenStudy (precal):

why don't you graph it and see what it looks like? You should focus on the critical parts of the quadratic. where it crosses the x axis, those locations are called the x intercepts where it crosses the y axis, that location is called the y intercept and the vertex because that is where the graph has symmetry ie x=1/4 is your axis of symmetry